OpenStudy (anonymous):

Find the equation of the curve which passes through the point (2,3) and for which dy/dx=3x^2+x.

OpenStudy (anonymous):

take the antiderivative of 3x^2+x, then find C so that it passes through (2,3)

OpenStudy (anonymous):

antiderivative or indefinite integral, I forget which

OpenStudy (anonymous):

Is it: dy/dx=3(2)^2+2=14 integrate dy/dx= x^3+x^2/2+c using pt (2,3), 3=2^3+2^2/2+c c=-3

OpenStudy (anonymous):

Therefore the eqn is: y=x^3+x^2/2-3

OpenStudy (anonymous):

Is it right?

OpenStudy (anonymous):

check finding c again...

OpenStudy (anonymous):

but I may be wrong.

OpenStudy (loser66):

c =-7 friend, so the equation is y = x^3 + x^2/2 -7